# Reflectionless and echo modes in asymmetric Damour-Solodukhin wormholes

### v2

@@ I Introduction
Pioneered by Nollert and Price [1, 2] and Aguirregabiria and Vishveshwara [3, 4], the black hole spectral instability has aroused much attention in the recent literature. The essence of such instability resides in the observation that insignificant modifications to the black hole metric, such as an approximation of the Regge-Wheeler effective potential in terms of piecewise functions [1, 5] or the introduction of some rather moderate discontinuities [6], will drastically deform the quasinormal mode [7, 8, 9] (QNM) spectrum, particularly the asymptotic behavior of high overtones [6]. This finding undermines the intuitive assertion that an insignificant modification of the effective potential shall not introduce a sizable impact on the resulting QNMs and has been attributed to the non-Hermitian nature [10, 11] of the gravitational system [12].
@@
The significance of spectral instability resides in its observational implications, especially regarding black hole spectroscopy [19]. The relevance of the topic is closely connected with the detections of gravitational waves emanating from the binary mergers by the LIGO and Virgo collaboration [20, 21, 22, 23]. The success of the ground-based facilities has further promoted the ongoing spaceborne projects, such as LISA [24], TianQin [25, 26], and Taiji [27], leading to speculation that direct observation of ringdown waveforms might also be plausible [28]. Under realistic astrophysical conditions, the source of gravitational waves is not a single isolated celestial object. Typically, they are submerged and interacting with surrounding matter. This leads to deviations from the ideal metric with perfect symmetries, motivating the studies of the so-called “dirty” black holes [29, 30, 31, 32] in the context of black hole perturbation theory. The obtained results on spectral stability warrant further scrutiny of these dirty compact objects. Regarding the deformed spectrum in the frequency domain, the asymptotic QNMs that align almost parallel to the real axis are closely related to the intriguing concept of echoes, a late-time phenomenon first speculated by Cardoso et al. [15, 33]. As potential observables, echoes were proposed to help distinguish different but otherwise similar gravitational systems via their distinct properties near the horizon. Such a physical picture has motivated many studies into echoes across various systems, including exotic compact objects such as gravastars [34, 35], boson stars [36], and particularly wormholes [37, 38, 39, 40, 41, 42, 43, 44, 45]. Similar to the late-time tail, echoes are also attributed to the analytic properties of Green’s function, as analyzed by Mark et al. [46]. In studies of Damour-Solodukhin type wormholes [47], Bueno et al. [48] have investigated echoes by explicitly solving for specific frequencies at which the transfer matrix becomes singular, providing further insights into the complex interplay of spacetime geometry and QNMs. It has been argued [49] that the spectral instability is closely associated with the black hole echoes, based on the frequency-domain analyses [16, 50, 51].
@@
More recently, the notion of quasi-reflectionless scattering modes (quasi-RSMs) was introduced in Ref. [83]. This proposal is motivated by the connection between greybody factors and ringdown waveforms. For symmetric cavity potentials, which mimic wormhole-like spacetime configurations, perfect RSMs exist at discrete real frequencies, whereas in more general situations the minima of the reflectivity correspond to quasi-RSMs. As an intriguing feature of the greybody factor in the high-frequency regime, these quasi-RSMs were shown to be primarily responsible for the echo phenomenon of the underlying compact object. These modes were found to be approximately uniformly distributed along the frequency axis.
Intriguingly, the above properties are reminiscent of those of echo modes discussed in [16, 51, 49], even though their definitions are rather different. The present study is therefore motivated to explore the underlying similarity between quasi-RSMs and echo modes. By promoting the definition of quasi-RSMs to fully reflectionless modes, these frequencies become complex and correspond to the poles of an appropriately modified Green’s function. In addition, we investigate asymmetric cavity potentials, specifically asymmetric Damour-Solodukhin wormholes, and demonstrate that the resemblance in the asymptotic properties of the reflectionless and echo-mode spectra can be largely obtained analytically. We carry out explicit derivations of the spectra of RSMs and echo modes using two different approaches, based on the evaluation of the scattering amplitudes and on the poles of the modified Green’s function. In general, the RSMs are found to lie closer to the real frequency axis, leading to more pronounced waveforms associated with RSMs than those related to echo modes for a given identical source.
@@ III Reflectionless scattering modes and their asymptotic properties
$U\equiv U(x_{c})=\begin{pmatrix}e^{-i\omega x_{c}}&0\\ 0&e^{i\omega x_{c}}\end{pmatrix}$ (25)
@@
$\mathbb{T}_{12}$ $=$ $e^{-2i\omega x_{c}}T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{2i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}=e^{-2i\omega x_{c}}\left[T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}\right],$ (27)
@@
$\mathbb{T}_{11}$ $=$ $T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{22}^{-1}+e^{4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{12}^{-1},$ (28)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}=0.$ (29)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{4i\omega x_{c}}\frac{T^{\mathrm{R}}_{12}}{T^{\mathrm{L}}_{12}}\frac{T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{11}}\frac{\left(T^{\mathrm{L}}\right)_{12}^{-1}}{\left(T^{\mathrm{L}}\right)_{21}^{-1}}\right]=0,$ (30)
@@
$\mathbb{T}_{22}$ $=$ $e^{-4i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}+T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{11}^{-1}$ (36)
@@
$\mathbb{T}_{21}$ $=$ $e^{-2i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{12}^{-1}+e^{2i\omega x_{c}}T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{22}^{-1}.$ (37)
@@
$e^{-4i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}+T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{11}^{-1}=0,$ (38)
@@
$T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{4i\omega x_{c}}\frac{T^{\mathrm{R}}_{22}T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{21}T^{\mathrm{L}}_{21}}\right]=0.$ (39)
@@ V.1 The spectra of quasi-RSMs, RSMs, and echo modes
$T=\begin{pmatrix}1-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\\ -i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&1+i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\end{pmatrix}$, (60)
@@
$T^{-1}=\begin{pmatrix}1+i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\\ i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&1-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\end{pmatrix}$. (61)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{4i\omega x_{c}}\frac{T^{\mathrm{R}}_{12}}{T^{\mathrm{L}}_{12}}\frac{T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{11}}\frac{\left(T^{\mathrm{L}}\right)_{12}^{-1}}{\left(T^{\mathrm{L}}\right)_{21}^{-1}}\right]=\left(1-i\frac{V_{0}^{\mathrm{R}}}{2\omega}\right)i\frac{V_{0}^{\mathrm{L}}}{2\omega}\left[1+e^{4i\omega x_{c}}\frac{V_{0}^{\mathrm{R}}}{V_{0}^{\mathrm{L}}}\frac{1+i\frac{V_{0}^{\mathrm{L}}}{2\omega}}{1-i\frac{V_{0}^{\mathrm{R}}}{2\omega}}\right]=0.$ (62)
@@
$\frac{V_{0}^{\mathrm{L}}}{V_{0}^{\mathrm{R}}}+e^{4i\omega x_{c}}=0,$ (63)
@@
$T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{4i\omega x_{c}}\frac{T^{\mathrm{R}}_{22}T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{21}T^{\mathrm{L}}_{21}}\right]=-\frac{V_{0}^{\mathrm{R}}}{2\omega}\frac{V_{0}^{\mathrm{L}}}{2\omega}-e^{4i\omega x_{c}}{\left(1+i\frac{V_{0}^{\mathrm{R}}}{2\omega}\right)\left(1+i\frac{V_{0}^{\mathrm{L}}}{2\omega}\right)}=0.$ (66)
@@
$\frac{V_{0}^{\mathrm{R}}V_{0}^{\mathrm{L}}}{4\omega^{2}}+e^{4i\omega x_{c}}=0,$ (67)
@@
${T}_{12}$ $=$ $i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
${T}_{21}$ $=$ $-i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
@@
$\left(T^{-1}\right)_{12}$ $=$ $-i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
$\left(T^{-1}\right)_{21}$ $=$ $i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$

### v3

@@ I Introduction
Pioneered by Nollert and Price [1, 2] and Aguirregabiria and Vishveshwara [3, 4], the black hole spectral instability has aroused much attention in the recent literature. The essence of such instability resides in the observation that insignificant modifications to the black hole metric, such as an approximation of the Regge-Wheeler effective potential in terms of piecewise functions [1, 5] or the introduction of some rather moderate discontinuities, will drastically deform the quasinormal mode [6, 7, 8] (QNM) spectrum, particularly the asymptotic behavior of high overtones [9]. This finding undermines the intuitive assertion that an insignificant modification of the effective potential shall not introduce a sizable impact on the resulting QNMs and has been attributed to the non-Hermitian nature [10, 11] of the gravitational system [12].
@@
The significance of spectral instability resides in its observational implications, especially regarding black hole spectroscopy [19]. The relevance of the topic is closely connected with the detections of gravitational waves emanating from the binary mergers by the LIGO and Virgo collaboration [20, 21, 22, 23]. The success of the ground-based facilities has further promoted the ongoing spaceborne projects, such as LISA [24], TianQin [25, 26], and Taiji [27], leading to speculation that direct observation of ringdown waveforms might also be plausible [28]. Under realistic astrophysical conditions, the source of gravitational waves is not a single isolated celestial object. Typically, they are submerged and interacting with surrounding matter. This leads to deviations from the ideal metric with perfect symmetries, motivating the studies of the so-called “dirty” black holes [29, 30, 31, 32] in the context of black hole perturbation theory. The obtained results on spectral stability warrant further scrutiny of these dirty compact objects. Regarding the deformed spectrum in the frequency domain, the asymptotic QNMs that align almost parallel to the real axis are closely related to the intriguing concept of echoes, a late-time phenomenon first speculated by Cardoso et al. [15, 33]. As potential observables, echoes were proposed to help distinguish different but otherwise similar gravitational systems via their distinct properties near the horizon. Such a physical picture has motivated many studies into echoes across various systems, including exotic compact objects such as gravastars [34, 35], boson stars [36], and particularly wormholes [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. Similar to the late-time tail, echoes are also attributed to the analytic properties of Green’s function, as analyzed by Mark et al. [50]. In studies of Damour-Solodukhin type wormholes [51], Bueno et al. [52] have investigated echoes by explicitly solving for specific frequencies at which the transfer matrix becomes singular, providing further insights into the complex interplay of spacetime geometry and QNMs. It has been argued [53] that the spectral instability is closely associated with the black hole echoes, based on the frequency-domain analyses [16, 54, 55].
@@
More recently, the notion of quasi-reflectionless scattering modes (quasi-RSMs) was introduced in Ref. [87]. This proposal is motivated by the connection between greybody factors and ringdown waveforms. For symmetric cavity potentials, which mimic wormhole-like spacetime configurations, perfect RSMs exist at discrete real frequencies, whereas in more general situations, the minima of the reflectivity correspond to quasi-RSMs. As an intriguing feature of the greybody factor in the high-frequency regime, these quasi-RSMs were shown to be primarily responsible for the echo phenomenon of the underlying compact object. These modes were found to be approximately uniformly distributed along the real frequency axis.
Intriguingly, the above properties are reminiscent of those of echo modes discussed in [16, 55, 53], even though their definitions are rather different. The present study is therefore motivated to explore the underlying similarity between quasi-RSMs and echo modes. By promoting the definition of quasi-RSMs to fully reflectionless modes [^1] After completing this study, we became aware of the work by Tuncer et al. [88], in which the notion of total absorption or virtual absorption mode is developed in the context of complex frequencies. From a physical standpoint, the definition proposed in [88] is essentially equivalent to the one adopted here, while the present work focuses on exploring these modes by explicitly incorporating the boundary conditions and the Green’s function., these frequencies become complex and correspond to the poles of an appropriately modified Green’s function. In addition, we investigate asymmetric cavity potentials, specifically asymmetric Damour-Solodukhin wormholes, and demonstrate that the resemblance in the asymptotic properties of the reflectionless and echo-mode spectra can be largely obtained analytically. We carry out explicit derivations of the spectra of RSMs and echo modes using two different approaches, based on the evaluation of the scattering amplitudes and on the poles of the modified Green’s function. In general, the RSMs are found to lie closer to the real frequency axis, leading to more pronounced waveforms associated with RSMs than those related to echo modes for a given identical source.
@@ III Reflectionless scattering modes and their asymptotic properties
$U\equiv U(x_{c})=\begin{pmatrix}e^{+i\omega x_{c}}&0\\ 0&e^{-i\omega x_{c}}\end{pmatrix}$ (25)
@@
$\mathbb{T}_{12}$ $=$ $e^{2i\omega x_{c}}T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{-2i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}=e^{2i\omega x_{c}}\left[T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{-4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}\right],$ (27)
@@
$\mathbb{T}_{11}$ $=$ $T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{22}^{-1}+e^{-4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{12}^{-1},$ (28)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}+e^{-4i\omega x_{c}}T^{\mathrm{R}}_{12}\left(T^{\mathrm{L}}\right)_{11}^{-1}=0.$ (29)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{-4i\omega x_{c}}\frac{T^{\mathrm{R}}_{12}}{T^{\mathrm{L}}_{12}}\frac{T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{11}}\frac{\left(T^{\mathrm{L}}\right)_{12}^{-1}}{\left(T^{\mathrm{L}}\right)_{21}^{-1}}\right]=0,$ (30)
@@
$\mathbb{T}_{22}$ $=$ $e^{4i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}+T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{11}^{-1}$ (36)
@@
$\mathbb{T}_{21}$ $=$ $e^{2i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{12}^{-1}+e^{-2i\omega x_{c}}T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{22}^{-1}.$ (37)
@@
$e^{4i\omega x_{c}}T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}+T^{\mathrm{R}}_{22}\left(T^{\mathrm{L}}\right)_{11}^{-1}=0,$ (38)
@@
$T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{-4i\omega x_{c}}\frac{T^{\mathrm{R}}_{22}T^{\mathrm{L}}_{11}}{T^{\mathrm{R}}_{21}T^{\mathrm{L}}_{21}}\right]=0.$ (39)
@@ V.1 The spectra of quasi-RSMs, RSMs, and echo modes
$T=\begin{pmatrix}1-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\\ i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&1+i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\end{pmatrix}$, (60)
@@
$T^{-1}=\begin{pmatrix}1+i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\\ -i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}&1-i\frac{V_{0}^{\mathrm{L,R}}}{2\omega}\end{pmatrix}$. (61)
@@
$T^{\mathrm{R}}_{11}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{-4i\omega x_{c}}\frac{T^{\mathrm{R}}_{12}}{T^{\mathrm{L}}_{12}}\frac{T^{\mathrm{L}}_{22}}{T^{\mathrm{R}}_{11}}\frac{\left(T^{\mathrm{L}}\right)_{12}^{-1}}{\left(T^{\mathrm{L}}\right)_{21}^{-1}}\right]=-\left(1-i\frac{V_{0}^{\mathrm{R}}}{2\omega}\right)i\frac{V_{0}^{\mathrm{L}}}{2\omega}\left[1+e^{-4i\omega x_{c}}\frac{V_{0}^{\mathrm{R}}}{V_{0}^{\mathrm{L}}}\frac{1+i\frac{V_{0}^{\mathrm{L}}}{2\omega}}{1-i\frac{V_{0}^{\mathrm{R}}}{2\omega}}\right]=0.$ (62)
@@
$\frac{V_{0}^{\mathrm{L}}}{V_{0}^{\mathrm{R}}}+e^{-4i\omega x_{c}}=0,$ (63)
@@
$T^{\mathrm{R}}_{21}\left(T^{\mathrm{L}}\right)_{21}^{-1}\left[1-e^{-4i\omega x_{c}}\frac{T^{\mathrm{R}}_{22}T^{\mathrm{L}}_{11}}{T^{\mathrm{R}}_{21}T^{\mathrm{L}}_{21}}\right]=\frac{V_{0}^{\mathrm{R}}}{2\omega}\frac{V_{0}^{\mathrm{L}}}{2\omega}+e^{-4i\omega x_{c}}{\left(1+i\frac{V_{0}^{\mathrm{R}}}{2\omega}\right)\left(1+i\frac{V_{0}^{\mathrm{L}}}{2\omega}\right)}=0.$ (66)
@@
$\frac{V_{0}^{\mathrm{R}}V_{0}^{\mathrm{L}}}{4\omega^{2}}+e^{-4i\omega x_{c}}=0,$ (67)
@@
${T}_{12}$ $=$ $-i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
${T}_{21}$ $=$ $i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
@@
$\left(T^{-1}\right)_{12}$ $=$ $i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$
$\left(T^{-1}\right)_{21}$ $=$ $-i\frac{1}{2}\frac{V^{\mathrm{L,R}}\sin\left(W^{\mathrm{L,R}}\sqrt{\omega^{2}-V^{\mathrm{L,R}}}\right)}{\omega\sqrt{\omega^{2}-V^{\mathrm{L,R}}}},$